You're riding a unicorn at 25m/s and come to a uniform stop at a red light 20m away. What's your acceleration ?
1
Expert's answer
2018-02-26T09:12:07-0500
0^2-v^2=2as
The acceleration is
a=-v^2/2s=-25^2/2(20) =-15.625 m/s^2 .
Answer: -15.625 m/s^2 .